OpenStudy (anonymous):

2.8 is what percent of 4?

OpenStudy (stamp):

\[4(x\%)=2.8\]\[4(\frac{x}{100})=2.8\]\[\frac{4x}{100}=2.8\]\[4x=2.8(100)\]\[x\ =\ ?\ \%\]

OpenStudy (stamp):

Alternative setup:\[\frac{2.8}{4}=\frac{x}{100}\]Cross multiply and solve for x.